OpenStudy (anguyennn):

How to find x intercept for x^4-4x^2?

OpenStudy (anonymous):

Set the y=0 So it would be \[0=x^4-4x^2\] Distribute it \[0=x^2(x^2-4)\] = \[0=x^2(x+2)(x-2)\]

OpenStudy (anguyennn):

so would x be -2, 2 and 0?

OpenStudy (anguyennn):

does the x^2 count

OpenStudy (anonymous):

Yes it does. Yep, you got it all correct.

OpenStudy (anguyennn):

oh and then what about 3x^2-x-2?

OpenStudy (anonymous):

\[3x^2-x-2=(3x+2)(x-1)\]

OpenStudy (anguyennn):

oh!!!!!

OpenStudy (anguyennn):

wow thank you so much

OpenStudy (anonymous):

No problem! Good luck on the rest.

OpenStudy (anguyennn):

I appreciate it :)

OpenStudy (anguyennn):

yes thank you.